Identify the number that is equal to five times the sum of the digits.
1
Expert's answer
2013-04-17T09:11:00-0400
1) Assuming that our two-digit number is ( x + y*10) , where y -the first digit , x - the second digit. And we know that (x+y)*5 must be equal to x + y*10.

2)So we have equation:
(x+y)*5 = x + 10y
5x + 5y = x + 10y
5x = x +10y - 5y
5x = x + 5y
5x - x = 5y
4x = 5y
0.8x = y

3) All digits must be integer , that's why x=5 , because it is the only digit that can make y integer. And we have y= 0.8*5 = 4.
Let's check the answer (4+5)*5 = 45 ALL RIGHT.

ANSWER IS 45